### v2

@@ Abstract
Inexact proximal augmented Lagrangian methods (pALMs) are particularly appealing for tackling convex constrained optimization problems because of their elegant convergence properties and strong practical performance. To solve the associated pALM subproblems, efficient methods such as Newton-type methods are essential. Consequently, the effectiveness of the inexact pALM hinges on the error criteria used to control the inexactness when solving these subproblems. However, existing inexact pALMs either rely on absolute-type error criteria (which may complicate implementation by necessitating the pre-specification of an infinite sequence of error tolerance parameters) or require an additional correction step to guarantee convergence when using relative-type error criteria (which can potentially degrade the practical performance of pALM). To address these deficiencies, this paper proposes ripALM, a relative-type inexact pALM, for linearly constrained convex optimization. This method can simplify practical implementation while preserving the appealing convergence properties of the classical absolute-type inexact pALM. To the best of our knowledge, ripALM is the first relative-type inexact version of the vanilla pALM with provable convergence guarantees. Numerical experiments on quadratically regularized optimal transport problems and basis pursuit denoising problems demonstrate the competitive efficiency of the proposed method compared to existing state-of-the-art methods. As our model encompasses a wide range of application problems, the proposed ripALM offers broad applicability and has the potential to serve as a basic optimization toolbox.
Keywords: augmented Lagrangian method; proximal term; relative-type error criterion; asymptotic (super)linear convergence rate; quadratically regularized optimal transport; basis pursuit denoising
@@ 1 Introduction
Constrained optimization, admitting excellent modeling power for real-world applications across a wide range of fields, including machine learning and data science, engineering, operations research, is a central area in optimization, especially in the current big-data era [7, 10]. This paper is dedicated to solving the following linearly constrained convex optimization problem:
@@
where $f:\mathbb{R}^{N}\to\mathbb{R}\cup\{+\infty\}$ is a (possibly nonsmooth) proper closed convex function, $A\in\mathbb{R}^{M\times N}$ and $\bm{b}\in\mathbb{R}^{M}$ are given data. Problem (1.1) is a classical and widely studied template in convex optimization. Moreover, since the smoothness of the objective function $f$ is not required in this work, problem (1.1) actually encompasses a much broader class of application problems than it may appear at first glance. For instance, by leveraging an indicator function of a closed convex set, problem (1.1) naturally accommodates more general convex constraints; see also two illustrative examples in Section 4.
The augmented Lagrangian method (ALM) is recognized as one of the most popular and effective methods for solving constrained optimization problems [19, 37]. The essential component of the ALM for solving problem (1.1) involves penalizing the linear constraint $A\bm{x}=\bm{b}$ to derive the augmented Lagrangian function, defined as
$\mathcal{L}_{\sigma}^{\rm prim}(\bm{x},\bm{y})=f(\bm{x})+\left\langle\bm{y},A\bm{x}-\bm{b}\right\rangle+\frac{\sigma}{2}\left\lVert A\bm{x}-\bm{b}\right\rVert^{2},\quad(\bm{x},\bm{y})\in\mathbb{R}^{N}\times\mathbb{R}^{M},$
where $\bm{y}\in\mathbb{R}^{M}$ is the Lagrangian multiplier associated with the linear constraint and $\sigma>0$ is a penalty parameter. Then, for a given sequence of penalty parameters $\{\sigma_{k}\}\subseteq\mathbb{R}_{++}$ and an initial Lagrangian multiplier $\bm{y}^{0}\in\mathbb{R}^{M}$, the ALM iteratively performs the following steps:
$\begin{cases}\bm{x}^{k+1}\in\arg\min\limits_{\bm{x}\in\mathbb{R}^{N}}\left\{\mathcal{L}_{\sigma_{k}}^{\rm prim}(\bm{x},\bm{y}^{k})\right\},\\[5.0pt] \bm{y}^{k+1}=\bm{y}^{k}+\sigma_{k}\nabla_{\bm{y}}\mathcal{L}_{\sigma_{k}}^{\rm prim}(\bm{x}^{k+1},\bm{y}^{k})=\bm{y}^{k}+\sigma_{k}\left(A\bm{x}^{k+1}-\bm{b}\right),\end{cases}$
which consists of minimizing one ALM-subproblem to get $\bm{x}^{k+1}$ and one gradient ascent step (associated with $\mathcal{L}_{\sigma_{k}}^{\rm prim}$) with a step size $\sigma_{k}$ to get $\bm{y}^{k+1}$. Since the augmented Lagrangian function $\mathcal{L}_{\sigma}^{\rm prim}$ is associated with the primal problem (1.1), the above ALM is usually named as a primal-based ALM. Alternatively, one may also consider a dual-based ALM which is associated with the dual problem of (1.1) given by (modulo a minus sign)
@@
where $f^{*}:\mathbb{R}^{N}\to\mathbb{R}\cup\{+\infty\}$ denotes the conjugate function of $f$ and $A^{\top}$ is the transpose of $A$. Then, given a penalty parameter $\sigma>0$, the augmented Lagrangian function associated with problem (1.2) is given by (see Section 2 for the detailed derivation)
$\mathcal{L}_{\sigma}^{\rm dual}(\bm{y},\bm{x})=-\bm{b}^{\top}\bm{y}+\frac{1}{2\sigma}\big\|\bm{x}+\sigma{A}^{\top}\bm{y}\big\|^{2}-\frac{1}{2\sigma}\|\bm{x}\|^{2}-\mathtt{M}_{\sigma f}\big(\bm{x}+\sigma{A}^{\top}\bm{y}\big),~~(\bm{x},\bm{y})\in\mathbb{R}^{N}\times\mathbb{R}^{M},$
where $\mathtt{M}_{\sigma f}(\bm{x}):=\min\limits_{\bm{z}}\left\{f(\bm{z})+\frac{1}{2\sigma}\|\bm{z}-\bm{x}\|^{2}\right\}$ denotes the Moreau envelope of the function $\sigma f$ at $\bm{x}$. Consequently, for a given sequence of penalty parameters $\{\sigma_{k}\}\subseteq\mathbb{R}_{++}$ and an initial point $\bm{x}^{0}\in\mathbb{R}^{N}$, the dual-based ALM reads as follows:
@@
The choice between using a primal-based or dual-based approach is problem-dependent. For example, when $f$ is nonsmooth, but its proximal mapping is easy-to-compute, the dual-based ALM may be more preferable. This is because, in the dual-based ALM, the optimization problem for updating $\bm{y}$ involves minimizing a continuously differentiable objective function, which can substantially simplify the optimization process.
The applicability of the ALM has been significantly expanded since Rockafellar’s seminal works [41, 42], which established a deep connection between the ALM and the proximal point algorithm (PPA). Moreover, the asymptotic (super)linear convergence rate of the ALM can be established under the assumption that the inverse of the subdifferential of the primal or dual objective function is Lipschitz continuous at the origin. Subsequent works have shown that this stringent condition can be further relaxed; see, for example, [11, 32] and references therein. Indeed, understanding the convergence properties of the ALM continues to be a topic of interest in the literature. Apart from theoretical advancements, numerous studies have demonstrated the remarkable efficiency of the ALM in solving various convex problems, including conic programming problems [23, 25, 27, 62], statistical optimization problems [22, 28, 60], optimization problems in machine learning [50, 54], and image/signal processing [21, 30], to mention just a few. These successful applications highlight the crucial roles of both the subsolvers employed to handle ALM subproblems and the error criteria used to control the accuracy required for solving these subproblems, while maintaining the desirable convergence properties of the ALM.
@@
To solve ALM-subproblems efficiently, second-order methods, such as Newton-type methods [38], are strong candidates due to their fast local convergence rates. For example, once the iterate of a Newton-type method falls into the fast convergence region, only a few more iterations are needed to achieve a highly accurate solution. However, certain regularity conditions are needed to ensure the nonsingularity of the (generalized) Hessian of the ALM subproblem’s objective function at its solution point. Theoretically, such regularity conditions are generally not verifiable in advance, except in some special cases [27]. Numerically, ill-conditioned or singular (generalized) Hessians can lead to unpredictable numerical behaviors, resulting in unstable implementations and excessive efforts for parameter tuning. To enhance the applicability of a second-order method for solving ALM’s subproblems, a common simple approach is to add a proximal term to the ALM subproblem’s objective function, resulting in the proximal augmented Lagrangian method (pALM) [^1] The pALM is also known as the proximal method of multipliers; see [41].. Specifically, given $\{\tau_{k}\}\subseteq\mathbb{R}_{++}$, $\{\sigma_{k}\}\subseteq\mathbb{R}_{++}$ and $(\bm{x}^{0},\bm{y}^{0})\in\mathbb{R}^{N}\times\mathbb{R}^{M}$, the dual-based pALM iteratively performs the following steps:
@@
$\begin{cases}\bm{y}^{k+1}=\arg\min\limits_{\bm{y}\in\mathbb{R}^{M}}\Big\{\mathcal{L}_{\sigma_{k}}^{\rm dual}(\bm{y},\bm{x}^{k})+\cfrac{\tau_{k}}{2\sigma_{k}}\left\|\bm{y}-\bm{y}^{k}\right\|^{2}\Big\},\\[8.0pt] \bm{x}^{k+1}=\bm{x}^{k}+\sigma_{k}\nabla_{\bm{x}}\mathcal{L}_{\sigma_{k}}^{\rm dual}(\bm{y}^{k+1},\bm{x}^{k})=\mathtt{prox}_{\sigma_{k}f}\big(\bm{x}^{k}+\sigma_{k}{A}^{\top}\bm{y}^{k+1}\big).\end{cases}$ (1.4)
By incorporating a proximal term, the pALM is able to bypass the regularity conditions when applying a second-order method for solving its subproblem. This typically results in more robust implementations with reduced parameter tuning. We refer readers to [12, 13, 18, 20, 24, 26, 28, 29, 36, 58] for successful applications of the pALM in addressing various important optimization problems.
To make the pALM truly implementable and practical, it must allow approximate solutions to the subproblem with progressively improved accuracy, and the associated error criterion must be practically verifiable, while preserving desirable convergence properties of the outer loop. To achieve this, an absolute-type error criterion was introduced by Rockafellar for the pALM in [41] and is now widely used in the literature; see, for example, [24, 26, 28, 29, 36]. The term “absolute” means that the sequence of error tolerances, which controls the accuracy required for solving each pALM subproblem, is predetermined and remains fixed throughout the iterations. However, such a sequence typically needs to be summable and must be carefully tuned to avoid being either overly conservative or excessively aggressive. Consequently, achieving desirable numerical performance often demands delicate hyperparameter tuning, which may incur considerable effort and potential inefficiencies in practical implementations. A feasible alternative is to adopt a relative-type error criterion. The term “relative” refers to controlling the errors in minimizing the subproblems based on certain computable quantities related to the current status of the algorithm. This relative-type criterion eliminates the need for a summable tolerance parameter sequence, thereby enhancing the practical implementability of the algorithm. The adoption of a relative-type error criterion started from the seminal works of Solodov and Svaiter [45, 46, 47, 48, 49] on inexact versions of PPA, and has since influenced the development of inexact versions of numerous algorithms, including the variable metric PPA [33], ALM [14], ADMM [15], FISTA [4], and the Bregman-type methods [56, 59]. However, to the best of our knowledge, research on relative-type inexact versions of the pALM remains in its early stage, with limited exploration in [13, 20, 58]. In particular, all these inexact versions are derived by applying a (variable metric) hybrid proximal extragradient method [33, 45, 46] to a certain primal-dual solution mapping of the convex constrained problem, and therefore require an additional correction step to ensure convergence. Consequently, the algorithms developed in [13, 20, 58] essentially deviate from the vanilla pALM framework (1.4). Moreover, our numerical results also indicate that the correction step tends to slow down the convergence of the pALM and thus increase the computational cost. These observations naturally lead us to raise the following question:
Can we design a relative-type error criterion for the vanilla pALM, while preserving desirable convergence properties?
In light of this, the primary goal of this paper is to address the above question by designing a relative-type error criterion for the vanilla pALM framework (1.4). Note that the relative-type error criterion proposed by Eckstein and Silva [14] for the ALM innovatively introduces an auxiliary error variable to control the inexactness while ensuring convergence. In this work, we shall demonstrate that a similar idea can be adapted to the vanilla pALM, but with essential modifications. We should emphasize that the corresponding theoretical convergence analysis also requires substantial modifications, due to the inclusion of the proximal term. Interestingly, while the initial motivation for considering pALM was to ensure the nonsingularity of the (generalized) Hessian of the ALM subproblem’s objective function at its solution point, our analysis reveals that this proximal term can further improve the existing convergence properties of the relative-type inexact ALM studied in [1, 14, 61]. Specifically, under standard assumptions, it ensures the convergence of both primal and dual sequences, as well as their asymptotic (super)linear convergence rates. The key contributions and findings of this paper are summarized as follows:
  • We develop a relative-type inexact proximal augmented Lagrangian method (ripALM) to solve the dual problem (1.2). To the best of our knowledge, the proposed ripALM is the first inexact version of the vanilla pALM (1.4). It not only shares the same convergence properties for the sequence $\{\bm{x}^{k}\}$ as the relative-type inexact ALM studied in [1, 14, 61], but also offers additional theoretical advantages: the sequence $\{\bm{y}^{k}\}$ is guaranteed to be bounded; moreover, under a relatively weaker [^2] As noted in [24, Lemma 2.4], the error bound condition in Assumption A used in this work is relatively weaker than the local upper Lipschitz continuity of $(\partial\ell)^{-1}$ at the origin. The latter condition was used in [61] to establish the asymptotic (super)linear convergence rate for the relative-type inexact ALM. For this reason, we use the term “weaker” here. error bound assumption, the whole sequence $\{\bm{y}^{k}\}$ is further proved to be convergent, and both primal and dual sequences achieve asymptotic (super)linear convergence rates. These results provide stronger theoretical support for the applicability of ripALM, particularly in scenarios where accurate solutions of both primal and dual problems are required.
  • Numerically, we implement the proposed ripALM and apply it to solve quadratically regularized optimal transport (QROT) problems and basis pursuit denoising (BPDN) problems. Extensive numerical results validate the promising performance of the proposed ripALM for solving large-scale QROT and BPDN problems. Moreover, comparisons with existing methods underscore our motivation and contributions in developing a relative-type error criterion for the vanilla pALM. In addition, since problem (1.1) has strong modeling capability, our ripALM can potentially be extended to solve optimal transport problems with other convex regularizers such as the group quadratic regularizer [58], provided that the regularizers are proximal-friendly and the generalized Jacobians associated with the pALM-subproblems are easy to compute. Given the growing interests in developing efficient algorithms for computational optimal transport [34], our work may open up new possibilities for applying the powerful pALM algorithmic framework to a wide variety of real-world applications.

### v3

@@ Abstract
Inexact proximal augmented Lagrangian methods (ipALMs) have been widely used for solving linearly constrained convex optimization problems, owing to their strong theoretical guarantees and excellent numerical performance. In practice, however, existing ipALMs typically employ Rockafellar-type absolute error criteria for solving the subproblems, which require delicate problem-dependent tuning of error-tolerance sequences. In this paper, we propose ripALM, a relative-type ipALM whose subproblem error criterion has only a single tolerance parameter in $[0,1)$. This makes the method simpler to implement and less sensitive to parameter tuning in practice. On the other hand, the use of such a relative-type error criterion renders the convergence of our ripALM beyond the scope of the convergence theory of existing ipALMs. To address this gap, we develop a new analysis framework under which ripALM is shown to admit desirable global convergence properties and it achieves an asymptotic (super)linear convergence rate under a standard error bound condition. While there exist other relative-type inexact pALMs, to ensure convergence, they require additional correct steps that generally impede the convergence speed. To the best of our knowledge, ripALM is the first relative-type inexact version of the vanilla pALM that avoids both summable tolerance parameter sequences and correction steps, while retaining rigorous convergence guarantees. Numerical experiments on quadratically regularized optimal transport and basis pursuit denoising problems demonstrate the effectiveness and robustness of our proposed method.
Keywords: proximal augmented Lagrangian method; relative-type error criterion; asymptotic (super-)linear convergence rate; quadratically regularized optimal transport; basis pursuit denoising
@@ 1 Introduction
Linearly constrained convex optimization is a central topic in optimization and arises in numerous applications in machine learning, data science, engineering, and operations research. In this paper, we consider the following classical model:
@@
where $f:\mathbb{R}^{N}\to\mathbb{R}\cup\{+\infty\}$ is a (possibly nonsmooth) proper closed convex function, $A\in\mathbb{R}^{M\times N}$ and $\bm{b}\in\mathbb{R}^{M}$ are given data. Notably, since the smoothness of the objective function $f$ is not required, problem (1.1) actually encompasses a broader class of application problems than it may appear at first glance. For instance, by incorporating indicator functions of suitable closed convex sets into the objective, problem (1.1) can accommodate more general convex constraints, provided that the resulting proximal mapping of the objective remains tractable; see two illustrative examples in Section 4.
Augmented Lagrangian methods (ALMs) and their variants are among the most widely used and effective methods for solving convex constrained optimization problems [14, 32]. They can be applied either to the primal problem (1.1) or to its associated dual problem, which, up to a sign change, can be written as
@@
where $f^{*}:\mathbb{R}^{N}\to\mathbb{R}\cup\{+\infty\}$ denotes the convex conjugate of $f$, and $A^{\top}$ is the transpose of $A$. The choice between primal-based and dual-based ALMs depends on the structure of the problem. In particular, when $f$ is nonsmooth but its proximal mapping is easy to compute, a dual-based approach is often preferable, because the corresponding ALM subproblem amounts to minimizing a continuously differentiable objective function and is therefore typically more tractable numerically. For this reason, we focus on the dual-based framework in this paper. Nevertheless, the analysis developed below applies equally to the primal-based setting.
More specifically, for a given penalty parameter $\sigma>0$, the augmented Lagrangian function associated with the dual problem (1.2) is given by
$\mathcal{L}_{\sigma}^{\rm dual}(\bm{y},\bm{x})=-\bm{b}^{\top}\bm{y}+\frac{1}{2\sigma}\big\|\bm{x}+\sigma A^{\top}\bm{y}\big\|^{2}-\frac{1}{2\sigma}\|\bm{x}\|^{2}-\mathtt{M}_{\sigma f}\big(\bm{x}+\sigma A^{\top}\bm{y}\big),\quad(\bm{x},\bm{y})\in\mathbb{R}^{N}\times\mathbb{R}^{M},$
where $\mathtt{M}_{\sigma f}(\bm{x}):=\min_{\bm{z}}\left\{f(\bm{z})+\frac{1}{2\sigma}\|\bm{z}-\bm{x}\|^{2}\right\}$ denotes the Moreau envelope of $\sigma f$ at $\bm{x}$; see Section 2 for the detailed derivation. Then, given a sequence of penalty parameters $\{\sigma_{k}\}\subseteq\mathbb{R}_{++}$ and an initial point $\bm{x}^{0}\in\mathbb{R}^{N}$, the dual-based ALM reads as follows:
@@
The applicability of the augmented Lagrangian method (ALM) has expanded substantially since Rockafellar’s seminal works [37, 38], which established its deep connection with the proximal point algorithm (PPA). This connection has led to a rich convergence theory for the ALM, including asymptotic (super)linear convergence under suitable error bound conditions; see, for example, [7, 27] and the references therein. Beyond its theoretical appeal, the ALM and its variants have demonstrated remarkable practical success in a wide range of applications, including conic programming [18, 20, 22, 52], statistical optimization [17, 23, 50], machine learning [43, 45], and image and signal processing [16, 25]. These developments also show that the effectiveness of the ALM depends critically not only on the outer augmented Lagrangian framework itself, but also on how its subproblems are solved and how the associated errors are controlled.
@@
To solve ALM subproblems efficiently, second-order methods, such as Newton-type methods [33], are particularly attractive because of their fast local convergence. In favorable regimes, only a few Newton-type iterations are often sufficient to obtain highly accurate subproblem solutions. However, their reliable performance typically depends on regularity conditions that ensure the nonsingularity of the (generalized) Hessian of the ALM subproblem objective at a solution. Such conditions are generally difficult to verify in advance and may even fail in practice, thereby limiting the applicability of second-order methods. A standard remedy is to add a proximal term to the ALM subproblem, which gives rise to the proximal augmented Lagrangian method (pALM) [^1] The pALM is also known as the proximal method of multipliers; see [37].. Specifically, given $\{\tau_{k}\}\subseteq\mathbb{R}_{++}$, $\{\sigma_{k}\}\subseteq\mathbb{R}_{++}$, and $(\bm{x}^{0},\bm{y}^{0})\in\mathbb{R}^{N}\times\mathbb{R}^{M}$, the dual-based pALM iteratively performs
@@
$\begin{cases}\bm{y}^{k+1}=\arg\min\limits_{\bm{y}\in\mathbb{R}^{M}}\Big\{\mathcal{L}_{\sigma_{k}}^{\rm dual}(\bm{y},\bm{x}^{k})+\cfrac{\tau_{k}}{2\sigma_{k}}\|\bm{y}-\bm{y}^{k}\|^{2}\Big\},\\[8.0pt] \bm{x}^{k+1}=\bm{x}^{k}+\sigma_{k}\nabla_{\bm{x}}\mathcal{L}_{\sigma_{k}}^{\rm dual}(\bm{y}^{k+1},\bm{x}^{k})=\mathtt{prox}_{\sigma_{k}f}\big(\bm{x}^{k}+\sigma_{k}A^{\top}\bm{y}^{k+1}\big).\end{cases}$ (1.4)
By incorporating this proximal term, the pALM improves the conditioning of the subproblem and facilitates the application of second-order methods as subsolvers. Consequently, it often exhibits more robust numerical performance in practice compared to the vanilla ALM. We refer the reader to [13, 19, 21, 23, 24, 31, 48] for successful applications of the pALM.
To make the pALM truly implementable and practical, one must allow its subproblems to be solved approximately under an error criterion that is both computationally verifiable and able to preserve desirable convergence properties of the outer loop. To this end, an absolute-type error criterion was introduced by Rockafellar for the pALM in [37] and has since been widely adopted in the literature; see, e.g., [19, 21, 23, 24, 31]. The term “absolute” means that the accuracy required for solving each pALM subproblem is controlled by a prescribed sequence of tolerance parameters. Such a sequence typically needs to be summable and must be chosen carefully to avoid being either overly conservative or excessively aggressive. Consequently, achieving satisfactory numerical performance often requires delicate parameter tuning, which can be laborious and complicate practical implementation.
An appealing alternative is to employ a relative-type error criterion. The term “relative” means that the errors in solving each subproblem are controlled by certain computable quantities related to the current progress of the algorithm, rather than by a prescribed sequence of tolerance parameters. Such a criterion often involves only a single tolerance parameter, and can therefore simplify parameter tuning and improve practical implementability. The use of relative-type error criteria can be traced back to the seminal works of Solodov and Svaiter [40, 41, 42] on relative-type inexact variants of PPA. Since then, this idea has influenced the development of relative-type inexact variants of numerous algorithms, including the variable metric PPA [28], ALM [9], ADMM [10], FISTA [4], and the Bregman-type methods [46, 49].
In contrast, relative-type inexact variants of the pALM remain much less explored, with only a few existing works; see [8, 15, 48]. More importantly, these existing methods are not genuine inexact versions of the vanilla pALM. They are all derived by applying a (variable metric) hybrid proximal extragradient method [28, 40, 41] to a suitable primal-dual solution mapping of the underlying convex constrained problem, and therefore require an additional correction step to guarantee convergence. As a result, the methods developed in [8, 15, 48] depart from the vanilla pALM (1.4). This distinction is not merely formal. Indeed, our numerical observations indicate that such a correction step tends to slow down practical convergence and increase the overall computational cost. It is therefore natural and important to ask:
Can one develop a relative-type inexact version of the vanilla pALM (1.4) that avoids both summable tolerance parameter sequences and correction steps, while still admitting rigorous convergence guarantees?
The primary objective of this paper is to answer the above question affirmatively. To this end, we propose a relative-type inexact proximal augmented Lagrangian method (ripALM) for solving the dual problem (1.2). The proposed method is a genuine inexact version of the vanilla pALM (1.4), whose subproblem error criterion has only a single tolerance parameter in $[0,1)$. This makes ripALM simpler to implement and more robust than existing absolute-type inexact pALMs, as observed from our numerical experiments in Section 5. However, the use of such a relative-type error criterion renders ripALM beyond the scope of the existing convergence theory for inexact pALMs. To overcome this difficulty, we develop a new analysis framework for ripALM. The main contributions of this paper are summarized as follows:
  • We introduce ripALM, which, to the best of our knowledge, is the first relative-type inexact version of the vanilla pALM (1.4) that avoids both summable tolerance parameter sequences and correction steps.
  • We develop a new analysis framework for ripALM, establish boundedness and global subsequential convergence properties of the generated iterates $\{\bm{y}^{k}\}$ and $\{\bm{x}^{k}\}$, and prove the convergence of the sequence $\{\bm{x}^{k}\}$. Moreover, under a standard error bound condition, we further prove the convergence of the sequence $\{\bm{y}^{k}\}$ and show that the sequence $\{(\bm{y}^{k},\bm{x}^{k})\}$ converges asymptotically at a (super)linear rate.
  • We demonstrate the practical effectiveness of ripALM on quadratically regularized optimal transport and basis pursuit denoising problems, where it is highly competitive with existing representative methods.